40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 240 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  5011   Thu Jul 21 10:08:30 2011 steveUpdateSUSsus sensor summary

OSEM voltages to be corrected at upcoming vent: threshold ~ 0.7-1.2V, ( at 22 out of 50 )

ITMX_UL, UR, LL, LR, SD

ITMY_UL

ETMX_UL, UR, LL, LR, SD

ETMY_UL

SRM_UL, UR, LL

MC2_UR_SD

BS_UR, SD

MC3_UL, LR, LL

 

Attachment 1: sus_sensor.png
sus_sensor.png
  5010   Thu Jul 21 09:04:59 2011 valera, steveUpdateSUSoplev gains were not optimized

 

Hi Steve,

 
I did change the ETMY optical lever configuration: http://131.215.115.52:8080/40m/4795
And I left it in that state per Jamie's request.
 
I was going to work on the servo tuning but found that the whitening was not working at that time.
What I was going to do is to measure the open loop gain to make sure the servo is stable, then 
measure the noise and minimize the rms motion by tuning the gain and the filter transfer function.
 
I plan to come to the 40m lab on August 22 for two weeks.
 
Valera.
  5009   Wed Jul 20 23:31:44 2011 SureshUpdateIOOMC Trans QPD is down

 

The mode cleaner is not locking because the MC Trans QPD signal is not present.  There is light on the QPD when the MC flashes and its position has not shifted.  The cable is plugged in well into the sensor head.  The signal cable is labled "MC2 Opt Lever"  and it arrives on the 1X4 rack along with other Optical Lever cables. Pressing the connector in did not solve the problem.

 

  5008   Wed Jul 20 22:16:27 2011 Ishwita, ManuelUpdateElectronicslying seismometer cable and plugging it

We laid the cable along the cable keeper from the BACARDI seismometer to the rack 1X6, the excess cable has been coiled under the X arm.

We plugged the cable to the seismometer and to the seismometer electronics box in rack 1X6. We also plugged the AC power cable from the box to an outlet in rack 1X7 (because the 1X6 outlets are full)

With the help of a function generator we tested the following labeled channels of AA board...

2, 3, 11, 12, 14, 15, 16, 18, 19, 20 and 24

that are the channels that can be viewed by the dataviewer, also the channel 10 can be viewed but it's labeld BAD so we cannot use it.

We leveled the seismometer and unlocked it, and saw his X,Y,Z velocity signals with an oscilloscope.

  5007   Wed Jul 20 20:44:56 2011 JamieUpdateSUSAll sus models rebuilt and restarted

There were a couple of recent improvements to the sus_single_control model that had not been propagated to all of the suspension controllers.

Rebuilt and restarted c1mcs, c1sus, c1scx, and c1scy.  Everything seems to be working fine after restart.

  5006   Wed Jul 20 20:04:54 2011 JamieUpdateCDSC1:DAQ-FB0_C1XXX_STATUS sometimes unexplainably goes red

I have been noticing this happening occasionally, but I don't understand what is causing:

status-fb-red1.png

The channel in question above is C1:DAQ-FB0_C1SCX_STATUS.  This channel is (I believe) reporting some status of the front end model communication with the frame builder, but I'm not sure exactly what.

Usually this problem goes away when I restart the model or the frame builder, but it didn't work this time.  Tomorrow I will figure out what this channel means, why it's sporadically going red, and how to correct it.

  5005   Wed Jul 20 19:48:03 2011 JamieUpdateSUSRe: oplev gains today

 

 We have been modifying models that need to have their channels renamed to run activateDQ when Joe's post_build_script to is run. 

The trick is to integrate things to get the post_build_script running after every model build (getting it hooked in to the make file somehow).  We're working on it.

 I've added the following epics channels to sus_single_control model using the epicsOutput part:

  • OL_SUM
  • OL_PITCH
  • OL_YAW

These channels are now all available.  I'm not exactly sure how to ensure that they're being trended.  I'll check that tomorrow.

  5004   Wed Jul 20 19:24:12 2011 ranaUpdateSUSoplev gains today

I guess Valera forgot to elog it. Steve, please email him and get the info.

I started to check out the OL servos today so that our whole interferometer is not too floppy.

  • The ETMX OPLEV DAQ channels were not in the list. Jamie ran the activateDQ.py script and it came back. Right now, we have no diagnostics to know if people have run this or not so the frames will have missing data now and again depending on how forgetful the rebooters are. Perhaps we can get activateDQ put into the make file???
  • I turned ON all of the offset buttons on the OL1, etc. filter banks. This allows for the dark offsets to be set for the OL quadrants. With these buttons off it doesn't make any sense.
  • I noticed that there are white (INVALID) fields all over the OPTLEV_SERVO screens. This is just because the new SUS models have not captured all of the functionality of the old system. Needs fixing.

Untitled.png

Some of these OL spectra are not like the others...

 a.png

  5003   Wed Jul 20 18:44:54 2011 KojiSummarySUSWeekly Summary

Find Frank and ask him about those components.

  5002   Wed Jul 20 17:43:33 2011 sureshUpdateComputersrestarted the frame builder

I restarted the frame builder in the last 15mins. 

I was making a change to a DAC channel in the C1IOO model.

  5001   Wed Jul 20 14:42:34 2011 steveUpdateSUSoplev gains today

C1:SUS-ETMX_OLPIT_GAIN set to 1.0     OLYAW 1.0

                ETMY                                     -0.2                   -0.2

                ITMY                                        2.0                    -2.0

                ITMX                                        0.5                     0.5

                  BS                                          0.4                    -0.4

                PRM                                         0.5                    -0.7

                SRM                                         1.0                     1.0

Earlier today Rana and I made power spectra of ETMY_OPLEV_ERROR signals with servo on and off.

It was indicating that the servo is not doing anything. These gain values were not set since IFO rebuilt.

Valera's entries were searched also. He did not do such thing. Rana may know where it is in the elog if it happened.

  5000   Wed Jul 20 12:05:08 2011 NicoleSummarySUSWeekly Summary

Since last week Wednesday, I have since found a Pomona Electronics box (thanks to Jenne)

to use for my photosensor head circuit (to house the LED and 2 photodiodes). Suresh has

shown me how to use the 9-pin Dsub connector punch, and I have punched a hole in this box

to attach the Dsub connector. 

 

Since this past entry regarding my mechanical design for the photosensor head (Photosensor Head Lessons),

I have modified the design to use a Teflon sheet instead of a copper PCB and I have moved the LED

and photodiodes closer together, upon the suggestions of Jamie and Koji.  The distance between

components is now 0.112" instead of the initial 0.28".  Last night, I cut the PCB board for the LED

and photodiodes and I drilled holes onto the PCB board and Teflon sheet so that the two may be

mounted to the metal plate face of the Pomona box.  I still need to cut the viewer hole for and

drill screws into the face plate.

P7200054.JPG

I have also been attempting to debug my photosensor circuit (box and LED/photodiode combination).

Since this last entry (Painful Votlage Regulator and Circuit Lessons), Suresh has helped me to get the parts

that I need from the Downs Electronics lab (15 wire ribbon cable, two 9 pin D-sub connectors M,

one 15 pin D-sub connector M, one 16 pin IDC connector). Upon the suggestion of Jamie, I have

also made additional safety changes to the circuit by fixing some of the soldering connections

so that all connections are done with wires (I had a few immediate lines connected with solder).

I believe the the photosensor circuit box is finally ready for testing. I may just need some help

attaching the IDC connector to the ribbon cable. After this, I would like to resume SAFELY

testing my circuit.

 P7200055.JPG

I have also been exploring SimMechanics. Unfortunately, I haven't been able to run the

inverted pendulum model by Sekiguchi Takanori. Everytime I attempt to run it, it says

there is an error and it shuts down Matlab. In the meanwhile, I have been watching

SimMechanics demos and trying to understand how to build a model. I'm thinking that

maybe once I figure out how SimMechanics works, I can use the image of his model

(I can see the model but it will not run) to construct a similar one that will hopefully work.

 

I have also been attempting to figure out the circuitry for the pre-assembled

accelerometer (made with the LIS3106AL chip).  I have been trying to use a multi-meter

to figure out what the components are (beyond the accelerometer chip, which I have

printed out the datasheet for), but have been unsuccessful at that. I have figured out

that the small 5 pin chip says LAMR and is a voltage regulator. I am hoping that if I can

find the data sheet for this voltage regulator, I can figure out the circuitry. Unfortunately,

I cannot find any datasheets for a LAMR voltage regulator. There is one by LAMAR, but

the ones I have seen are all much larger. Does anyone know what the miniature voltage

regulator below is called and if "LAMR" is short for "LAMAR"?

 

P7200056.JPG

 

  4999   Wed Jul 20 11:42:47 2011 Ishwita, ManuelUpdate Weekly summary
  • We gave a white-board presentation on derivation of formula for optimum Wiener filter coefficients and wrote a latex document for the same. relevant elog entry
  • We enjoyed drilling the cover of the AA board and fixing it.
  • The AA board was fixed on rack 1X7 with Jenne's help. relevant elog entry
  • We tried writing a simulation for the transfer function of the stacks in Matlab. Once we get some satisfying results, we will post it on the elog.
  • We started reading the book 'Digital Signal Processing - Alan V. Oppenheim / Ronald W. Schafer' and are still reading it. We also tried watching lecture videos on z-transform...
  4998   Wed Jul 20 11:13:59 2011 Larisa ThorneUpdateelogI restarted the ELOG as it seemed to have crashed

 

  4997   Wed Jul 20 10:10:19 2011 SonaliUpdateGreen LockingWeekly summary

 I finished wih the set-up at the ETMY table. Instead of the neutral Density Filter , I put in a mirror(Y1-1037-45S)  which is reflective for IR , so that only 1% of the light is incident on the fibre  as per Rana's suggestion.

Now, the power incident on the fibre is measured to be 6 mW and the power measured out of the fibre is 2.76 mW after the necessary alignments.

On the PSL able, I have routed the beam that is coming out of the back of the PMC(instead of the dumped light from the oven to prevent any light from reflecting back into the laser), to the area where I am putting the set-up for the superposition of the PSL and the ETMX and ETMY beams.

Today I will proceed with the layout.

  4996   Wed Jul 20 06:52:01 2011 SureshUpdateLSCABSL - PSL beat lock

The ABSL locking setup to the PSL is down. 

According the plan, I started to use the IR beam dumped after the doubling crystal for the IR beat lock (Sonali's project).  The beat lock was disturbed when I shifted some clamps to make way for a few mirrors.  So I set about fixing the beat lock.  I reobtained the lock but noticed that the net beam power reaching the Newfocus 1611 detector was around 15mW.  10mW from the ABSL and 5mW from PSL.

This is much too high as the maximum allowed on 1611 is 2mW. 

I therefore started to adjust the power levels by using  Y1-1064-45S mirrors at non-45 deg angles.  However Rana pointed out that this would lead to amplitude noise due to the mirror vibrations.  I then switched to using beam splitters as pick offs.   This is better than using neutral density filters since the back scatter is lower this way.

David wanted some of the ABSL beam for his SURF student.  So I changed the mirror after beam expanding telescope on the ABSL route to provide this power.  We also installed a pair of half wave plates and a PBS to allow us smooth power level control on this beam.

The beat lock setup is now down and needs to be completed for PRCL and SRCL measurements.

 

 

  4995   Wed Jul 20 06:36:39 2011 SureshUpdateIOOWFS1 and 2 gains

Gains of individual quadrants in both the WFS

As a simple check of the gains on all the quadrants I hooked up the AM (Jenne) laser to put FM modulated light on to the WFS heads and observed the FM modulation frequency , 105 Hz, show up on a power spectrum of the RF outputs.   The plots below show the peak at 105Hz in all the quadrants.

WFS1_7kHz_105Hz.pdf

WFS2_7kHz_105Hz.pdf

 

However I should have put in AM modulation rather FM modulation.  I will do that using the digital system today.  The first version above was done wth a Marconi driving the AM laser modulation.

 

 

  4994   Wed Jul 20 06:17:04 2011 SureshUpdateGreen LockingY-end green laser power issues

The Y-end green beam power is 0.47 mW.

While aligning the Y-end aux laser light into the fiber we noticed that the green power out of the doubling crystal was in microwatts.  I checked to see what was the trouble and found that the oven was cold as the temperature controller had been disabled.  I enabled it and scanned the temperature to maximise the green output.  Yet the power is less than 10% of that at the X end (7mW).

To verify I checked the power of various beams on the Y-end table.  They are listed below in the picture

Y-end_table_powers.pdf

The green beam power is proportional to the square of the IR incident power and this explains the drop in green power by a factor of (210/730)^2  thus making 7 mW -->  0.5 mW.  However we may be able to double the power at the Y-arm oven if the uncoated lenses in the IR path are exchaned for coated ones. 

 

The green beam injection into the Y-arm cavity also needs to be cleaned up as noted here.  As seen in the picture below two of the mirrors which launch the beam into the arm cavity need to be fixed as well.

 Y-end_table.png




  4993   Tue Jul 19 23:39:11 2011 Jamie, JenneUpdateLSCMajor overhaul of LSC rack; binary switching of whitening filters now working

Yesterday we started going through the LSC binary whitening switching to make sure the new switching control in the LSC model was working.  Jenne and I hooked up a fancy home-brew white noise generator [0] into all of the LSC whitening filter inputs and started switching the whitening filters to see what would happen.  We found that some of the channels were switching, but the majority were not, or worse yet switching the wrong channels.  Upon closer inspection, and after finally finding the LSC wiring schematic, we found that the LSC rack cross-connect/back-plane cabling was pretty much a complete mess, and didn't at all correspond to the channel layout in Suresh's diagram.

Given that the back-plane wiring had to be almost completely redone, we decided to completely redo the LSC electronics layout, to be a little more self-consistent and to use the given space more efficiently.  We'll post an updated electronics drawing soon.  The LSC model was also updated to reflect the new layout.

We then went through and verified that all of the whitening switching was working with the new layout.  As described previously, the first filter in the PD input filter bank is used to control the switching.  We did indeed verify that all the switching is working, but we noticed that switching logic was inverted, such that the whitening filter engaged when the filter was turned off.  This was fixed in the model and all the switching logic was verified to be working as expected.

Everything has now been hooked back up, but we need to verify that we're getting all of the PD demodulated RF and DC outputs as expected.  We need to check the RF phases, as some of the RF cable lengths have changed.

[0] a 50k resistor

Links:

 

  4992   Tue Jul 19 21:05:55 2011 haixingUpdateDAQchoose the right relay

Rana and I are working on the AA/AI circuit for Cymac. We need relays to bypass certain paths in the circuit, and we just found a nice website
explaining how to choose the right relay:

http:/zone.ni.com/devzone/cda/tut/p/id/2774

This piece of information could be useful for others.

  4991   Tue Jul 19 20:36:08 2011 ranaUpdateComputersVirtualBox + Windows 7 on rossa

I installed Virtual Box on rossa. Then I put Windows 7 in there and am now installing Altium.

You can run Windows on rossa by just clicking Applications -> System Tools -> Virtual Box.

  4990   Tue Jul 19 13:45:22 2011 SureshUpdateSAFETYPainful Voltage Regulator and Circuit-Testing Lessons

There is should be a few IDC connectors in the lab (and some ribbon cable) using which you can proceed with the testing of the circuit, if you prefer.  If not we can get them from our ever helpful electronics division at Downs.  In any case there is no need to lose time waiting for parts to arrive.

 

  4989   Tue Jul 19 10:54:14 2011 steveUpdateGenerallarge sensor card can not be found

Please return sensor card to  laser log box so others can use it. We have only one larger fluorescent sensor card.

  4988   Tue Jul 19 10:18:24 2011 ranaUpdateLSCBig ol' mess

Remember, as per our marker board conversation, that the resistor noise as excitation method only works if the gain of all of the channels is set to something high (like 45 dB).

At 0 dB, the resistor noise is only 30 nV/rHz, whereas the ADC noise is more like 10000 nV/rHz...

  4987   Tue Jul 19 09:19:14 2011 steveUpdateSUSPRM damping restored

The PRM watchdogs were tripped. The side was kicked up to 180mV Damping was restored.

  4986   Mon Jul 18 22:26:43 2011 NicoleUpdateSAFETYPainful Voltage Regulator and Circuit-Testing Lessons

Today I learned some important circuit-building lessons while testing my photosensor circuit box (i.e. how NOT to test a circuit and, conversely, things that should be done instead). 

I blew my first circuit today. The victim is in the photo below (bottom 7805 voltage regulator). The plastic covering fell off after I removed the fried regulator.  After checking various components, I figured out that I blew the circuit because I had forgotten to ground the regulator.  Although this was very unfortunate, I did make an important discovery. While testing the voltage output of the 7805 voltage regulator (I put a new one), I discovered that contrary to the claims of the datasheet, an input voltage of 5V will not produce a steady 5V supply. I found that at 5V, my regulator was only producing 4.117 V. I was using a 5 V supply because I wanted to use only 1 power supply (I was using a two-channel power supply that had a fixed 5V output to produce the +15, -15, ground, and 5 V I need for my photosensor circuit box).  After seeing this, I got a second power supply and am now using 10V to as an input for the regulator to produce 4.961V. I found that from a voltage range of 10V to 15 V, the regulator produced a steady  4.961 V supply. I have decided to use 10V as an input. My newly-grounded voltage regulator did not smoke or get hot at 10V.

After several more debugging trials (my LED was still not lighting up, according to the infared viewer), I learned another painful lesson. I learned DO NOT USE CLIP LEADS TO TEST CIRCUITS!!!! Initally, I was powering my circuit and making all of my connections between the photosensor head (2 photodiodes and 1 LED) with clip leads. This was a BAD IDEA BECAUSE CLIP LEADS ARE UNSTABLE AND IT IS VERY EASY TO SHORT A CIRCUIT IF THEY ACCIDENTALLY TOUCH! I did not realize this important lesson until my photosensor circuit was once again burning. Confused as to why my circuit was once again burning, I foolishly touched the voltage regulator. As you can see on the top voltage regulator in the photo below, my finger left its mark on the smoldering voltage regulator. As you cannot see the wincing on my face as I try to type this long elog, I will painfully type that the voltage regulator left its own mark on my finger (an ugly sore little welt).  Suresh has taught me a valuable lesson: WHEN DEALING WITH SOMETHING OF QUESTIONABLE/UNKNOWN TEMPERATURE, USE YOUR NOSE, NOT YOUR FINGER TO DETERMINE IF THAT COMPONENT IS HOT!!!! 

P7180424.JPG

 

To make my circuit-testing safer, upon the suggestion of Suresh, I have since removed the clip leads and inserted a 12 pin IDC component (pictured below). There are 12 pins for the 6 inputs I will get from each of the 2 photosensor heads. I have requested orders for a 16 pin IDC connector, 15 pin Dsub male part, 15 pin Dsub feed-thru, 9 pin Dsub male part (2), and 9 pin Dsub feed-thru (2). After receiving these components, I should be able to safely test my circuit.

P7180423.JPG

 In the meanwhile, I can explore SimMechanics and try to figure out how to use the accelerometer

  4985   Mon Jul 18 21:06:32 2011 PSL Table GuardianOmnistructurePSLDon't leave the PSL table open, unattended!!!!!!!!!!!!11111

I found the PSL table left open, and unattended again. 

As far as I know, Jamie and Jenne (working on the LSC rack, so no lasers / optics work involved) have been the only ones in the IFO room for several hours now. 

I'm going to start taking laser keys, or finding other suitable punishments.  Like a day of lab cleanup chores or something.  Seriously, don't leave the PSL table open if you're not actively working on it.

  4984   Mon Jul 18 20:59:19 2011 JenneUpdateLSCBig ol' mess

[Jamie, Jenne]

We decided to take on the deceptively easy-sounding task of checking that the LSC whitening switching was happening as anticipated.  We hoped to discover that when we clicked the "unwhitening" switches in FM1 of the LSC PDs, we would see the analog whitening turn on and off for the matching channel.  That is what is supposed to happen.

Tragically, it is instead one big giant crazy disaster of a mess.

What we did:

Made a 24tapus (octopus like last time, except more...), with a 50kOhm resistor as our white noise source (instead of using a DAC channel and AWG). 

We plugged our 24tapus into the 3 of 4 whitening boards on the LSC rack that are currently in use.  One of the boards just has 8 terminators on the input, so we left that one alone for now. 

We put the whitening gains to 0dB so that all the channels looked the same. 

We looked at the PD _IN1 channels in DTT, and monitored which signals had whitening switching when we clicked the "unwhitening" buttons on the PD filter banks. 

So far, we can find no rhyme or reason as to why some of the channels work (click unwhite on that PD, see that signal have whitening switching), and others don't.  Some channels we just can't get to switch no matter what, others are just mis-mapped.  There is no discernible pattern.

What we think (so far) is going on:

All of the cables from the PD demod boards are going to the Whitening board inputs, exactly as in Suresh's Diagram.  The only difference is that Refl33, AS165 and Refl165 demod boards don't exist in the rack at this time. 

The Whitening and AA boards in Suresh's Diagram labeled 0-7 are connected to Binary Output channels 0-7. This is a good thing.

The Whitening and AA boards in the diagram labeled 8-15 are connected to Binary Output channels 24-31. This is not so awesome.

This is all we are confident about at this time.

Next steps:

We are hoping that Ben has a secret stash (or can tell us who would) of LSC rack wiring diagrams.  We would like to find out, without the pain of tracing wires and cables by hand, how the Binary I/O information gets through the cross-connect on the LSC rack up to the whitening boards. 

We are leaving the 24tapus in place for now, so that we can carry on tomorrow, either with a wiring diagram in hand, or carefully tracing cables. 

  4983   Mon Jul 18 15:51:04 2011 steveUpdateSUSETMY spectra of oplev error signal
Attachment 1: ETMYopler.png
ETMYopler.png
  4982   Mon Jul 18 14:39:53 2011 JenneUpdateComputersLSC PD chans acquired

There's too much tromping around, so I'm not going to actually measure PRC length right now, but I did set some channels to be acquired (POPDC, POXDC, POYDC) in addition to ASDC which was already acquired, so that I can look at the resonance fringes when I sweep the ABSL laser (hopefully later tonight....)

  4981   Mon Jul 18 10:24:48 2011 steveHowToGeneralsolvent bottles for cleaning optics

Each bottle has matched seals. They are not interchangeable.

It is critical that the solvent do not reach the rubber bulb. Practice with the pipet.

In case of solvent touching the suction bulb: do not let the solvent go back into the bottle! Remove bulb, let it dry out and rinse pipet.

It is essential that the solvent bottle must be rinsed and refilled if it's content met with the rubber bulb.

Use glass syringe with SS needle in critical application: Hamilton ~0.1 ml

Attachment 1: P1080090.JPG
P1080090.JPG
  4980   Sun Jul 17 18:23:23 2011 JenneUpdatePSLFound the PMC unlocked

It was unlocked since ~4:30am.  No idea why.  It's relocked so I can try round N of measuring the PRC length.

Attachment 1: PMCunlocked_17July2011.png
PMCunlocked_17July2011.png
  4979   Sat Jul 16 18:54:05 2011 Ishwita, ManuelConfigurationElectronicsAA board

We fixed the anti-aliasing board in its aluminum black box,  the box couldn't be covered entirely because of the outgoing wires of the BNC connectors, so we drilled additional holes on the top cover to slide it backwards by 1cm and then screw it.

We had to fix the AA board box in rack 1X7, but there wasn't enough space, so we tried to move the blue chassis (ligo electro-optical fanout chassis 1X7) up with the help of a jack. We removed the blue chassis' screws but we couldn't move it up because of a piece of metal screwed above the blue chassis, then we weren't able to screw the two bottom screws again anymore because it had slided a bit down. Thus, the blue chassis (LIGO ELECTRO-OPTICAL FANOUT CHASSIS 1X7) is still not fixed properly and is sitting on the jack.

To accommodate the AA board (along with the panel-mounted BNC connectors) in rack 1X7 we removed the sliding tray (which was above the CPU) and fixed it there. Now the sliding tray is under the drill press.

 

Attachment 1: DSC_3236.JPG
DSC_3236.JPG
Attachment 2: pic1.png
pic1.png
Attachment 3: DSC_3237.JPG
DSC_3237.JPG
  4978   Fri Jul 15 19:00:18 2011 dmassMetaphysicselogCrashes

Elog crashed a couple times, restarted it a couple times.

  4977   Fri Jul 15 17:42:21 2011 SonaliUpdateGreen LockingPSL layout for superposition of the PSL,ETMX and ETMY beams.

The fibres carrying the beams from the ETMX as well as the ETMY have been routed to the PSL table now.

A part of the PSL beam has to be superposed on the fibre-outputs to obtain a beat signal. We have located a stray beam on the PSL(which is currently being dumped) which we plan to redirect for the same. The layout of the plan is attached herewith.

Attachment 1: PSL_change1.pdf
PSL_change1.pdf
Attachment 2: PSLnewspace1.pdf
PSLnewspace1.pdf
  4976   Fri Jul 15 16:14:00 2011 steveUpdateSUSBS oplev error signal spectra
Attachment 1: BS_opl_ersig.pdf
BS_opl_ersig.pdf
  4975   Fri Jul 15 14:29:30 2011 NicoleSummarySUSPhotosensor Head Lessons

 

 You are right Jamie! Thank you for the correction! I will now use the Teflon sheet instead of the PCB piece.

The photodiodes do have three legs, but I imagined the third one lying on a different plane, since it is spaced apart from the two I have drawn.

I should include this third leg in my drawing?

  4974   Fri Jul 15 14:23:30 2011 NicoleSummarySUSPhotosensor Head Lessons

 

 Ah! I see! Thank you!

I should put the LEDs and photodiodes closer together so that more of the reflected light falls on the photodiodes and the photodiodes have a higher response.

Also the reflectivity of the mirror will be optimized if the incident light is normal to the mirror surface. We will be setting up the photosensor and mirror so that the LEDs

emit light normal to the mirror surfaceDuring displacement, this light may be slightly off-normal but still close to normal incidence. We want the photodiodes to be close to the LED since we want

them to detect light that is close to the path of normal incidence (small angles of reflection). [Thanks to Jenne for helping me figure this one out!]

 

Thank you for the suggestion ^___^

  4973   Fri Jul 15 13:48:56 2011 sonaliUpdateGreen LockingPower of the AUX laser increased.

Quote:

Quote:

The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).

The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.

 For the phase locking or beat note measuring we only need ~1 mW. Its a bad idea to send so much power into the fiber because of SBS and safety. The power should be lowered until the output at the PSL is < 2 mW. In terms of SNR, there's no advantage to use such high powers.

 

Well,the plan is to put in  a neutral density filter in the beam path before it enters the fibre. But before I could do that, I set up the camera on the PSL table to look at the fiber output . I will need it while I realign the  beam after putting in the Neutral Density Filter. I have attached the ETMY layout with the Neutral Density filter in place herewith.

Attachment 1: ETMY_after_fibre_coupling_labelled.pdf
ETMY_after_fibre_coupling_labelled.pdf
  4972   Fri Jul 15 09:25:02 2011 ranaUpdateSUSSUS oplev spectras

In addition to the OL quadrants, you need to plot the OPLEV_PERROR and OPLEV_YERROR signals since these are the real signals we use for finding the mirror motion. If they're not in the Dataviewer, Jamie should add them as 256 Hz DAQ channels (using these names so that we have the continuity with the past). These DAQ channels correspond to the IN1 channels for the OL filter banks.

Also JPG are banned from the elog - you should put all of the plots into a single, multipage PDF file in honor of the new Wagonga.

  4971   Fri Jul 15 08:48:36 2011 JamieSummarySUSPhotosensor Head Lessons

Nicole: I thought we had decided to use teflon as the insulator between the PCB (yellow) and the LED/PDs?  I don't think you should use another circuit board with copper on it.  The copper will short the LED/PD heads to the metal box, which might be problematic.

Otherwise the design looks pretty good.  I think the PDs have three leads each, yes?

  4970   Fri Jul 15 01:11:21 2011 KojiSummarySUSPhotosensor Head Lessons

Rotate the PDs and the LED so that you can put them as close as possible.
This is to increase the sensitivity of the sensor. Think why the closer the better.

  4969   Thu Jul 14 20:24:32 2011 NicoleSummarySUSPhotosensor Head Lessons

Today I tested the photosensor head combination (2 Hamamatsu S5971 photodiodes and 1 Hamamatsu L9337 LED). I discovered that I had burnt out the LED and the photodiodes when I soldered them to the PCB board.

After looking up soldering information on Hamamatsu photodiodes, I learned that I need to solder at least 2 mm away from the head. I checked the pins of my burnt-out photodiodes and I had soldered 1.5 mm away from the head. To prevent this problem from happening again, Suresh suggested that I clip a lead onto photodiode/LED pin while I solder on connections to help dissipate some of the heat.

Today I was able to get a single photodiode (not attached to the PCB) to measure light emitted from an LED and I observed how voltage fluctuated as I moved the photodiode around the LED.

Suresh and Jamie also helped me to fix my photosensor head design (to make it more electrically-stable). Originally, I had planned to solder the LED and photodiodes onto a PCB and to mount that PCB to the front of a small metal Pomona Electronics box (with a whole cut out for the photodiodes and LED) using spacers, screws, and nuts.  However, the PCB I am using to solder on the LED and photodiodes has metal connections that may cause problems for the LED and photodiodes lying on the surface. Now, the plan is to have the LED and photodiodes mounted to the PCB with an insulatory PCB in between. Below is an explanatory picture.  I will determine the placement of the LED and photodiodes after making screws holes in the two PCBs to attach to the metal face of the box. I want to attach the screw holes first to make sure that the PCBs (and attached photosensor) are centered.

 

Photosensor_Head.jpg

  4968   Thu Jul 14 17:34:35 2011 Ishwita, ManuelHowToWienerFilteringWiener-Hopf equations

Since we are using Wiener filtering in our project, we studied the derivation of Wiener-Hopf equations. Whatever we understood we have written it as a pdf document which is attached below...

Attachment 1: derivwf.pdf
derivwf.pdf derivwf.pdf derivwf.pdf derivwf.pdf derivwf.pdf derivwf.pdf
  4967   Thu Jul 14 15:27:08 2011 steve,UpdateSUSSUS oplev spectras

Quote:

Quote:

Quote:

Healthy BS oplev

 I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948  The oplev servo was left ON!

 It is working today! Finally I repeated the BS spectra, that we did with Kiwamu last week

 The optical levers were centered during these measurements  without the reference of locked cavities.  They have no reference value now.

SRM sus need some help. ITMX is showing pitch/yaw modes of the pendulum .....OSEM damping is weak?

Attachment 1: BS_oplev.jpg
BS_oplev.jpg
Attachment 2: PRM_oplev.jpg
PRM_oplev.jpg
Attachment 3: ITMX_oplev.jpg
ITMX_oplev.jpg
Attachment 4: ETMX_oplev.jpg
ETMX_oplev.jpg
Attachment 5: ETMY_oplev.jpg
ETMY_oplev.jpg
Attachment 6: SRM_oplev.jpg
SRM_oplev.jpg
Attachment 7: ITMY_oplev_b.jpg
ITMY_oplev_b.jpg
  4966   Thu Jul 14 09:38:50 2011 steve,UpdateSUSBS oplev spectra

Quote:

Quote:

Healthy BS oplev

 I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948  The oplev servo was left ON!

 It is working today! Finally I repeated the BS spectra, that we did with Kiwamu last week

Attachment 1: BS_oplev.jpg
BS_oplev.jpg
  4965   Thu Jul 14 02:32:11 2011 sonaliUpdateGreen LockingPower of the AUX laser increased.

Quote:

The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).

The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.

 For the phase locking or beat note measuring we only need ~1 mW. Its a bad idea to send so much power into the fiber because of SBS and safety. The power should be lowered until the output at the PSL is < 2 mW. In terms of SNR, there's no advantage to use such high powers.

  4964   Wed Jul 13 12:24:46 2011 NicoleUpdateSUSWeekly Update
This week, I have been working on the photosensor circuit box.  This photosensor box will contain the current-stabilizing power supply and
voltage readout for the two photosensors I plan to build.
 
Suresh helped to walk me through the design of the photosensor circuit (image below) so I now understand how the circuit works.
PHOTOSENSORPLAN.JPG
 
 
Jaimie helped me to reorganize the original circuit layout I had to make it easier to follow. I have now redone half of the circuit (enough for one LED and photodiode pair). I still need to put in the voltage-regulators to provide the + and - 15 V neeCto power the op-amps but I will do that after testing the circuit.
prelimcircuit.JPG

In order to test this preliminary circuit, I need to build the photosensor heads.  Yesterday, Suresh helped me to open one of the professionally-build photosensors in the lab to understand how to arrange my photosensor heads. I now understand that I need to rigidly-mount the PCB to photosensor head box. I plan to use the PCB below. It will be sufficient for the lower-frequency range (below 10Hz) that I am interested in. 

PCBforphotosensor.JPG

 I would like to use a metal box like the one below to make each photosensor head. I looked in the lab last night for similar boxes but could not find one. Does anyone know where I can find a similar metal box?

lookingforbox.JPG

 

I am now working on accelerometer. I am working on attaching these metal wires to the pins of the accelerometer so that I can use clip leads to power and extract voltage measurements from my circuit.

 accelerometer.JPG

  4963   Tue Jul 12 17:30:24 2011 steve,UpdateSUSBS oplev spectra

Quote:

Healthy BS oplev

 I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948  The oplev servo was left ON!

Attachment 1: BS_oplev.jpg
BS_oplev.jpg
  4962   Tue Jul 12 11:52:54 2011 Jamie, SureshUpdateLSCLSC model updates

The LSC model has been updated:

Binary outputs to control whitening filter switching

We now take the filter state bit from the first filter bank in all RF PD I/Q filter banks (AS55_I, REFL11_Q, etc) as the controls for the binary analog whitening switching on the RF PD I/Q inputs. The RF_PD part was also modified to output this control bit. The bits from the individual PDs are then combined into the various words that are written to the Contec BO part.

Channel mapping updated/fixed to reflect wiring specification

Yesterday Suresh posted an updated LSC wiring diagram, with correct channel assignments for the RF PD I/Q and DC inputs.  Upon inspection of the physical hardware we found that some of LSC the wiring was incorrect, with I/Q channels swapped, and some of the PDs in the wrong channels.  We went through and fixed the physical wiring to reflect the diagram.  This almost certainly will affect the EPICS settings for some of the input channels, such as offsets and RD rotations.  We should therefore go through all of the RF inputs and make sure everything is kosher.

I also fixed all of the wiring in the LSC model to also reflect the diagram.

Once this was all done, I rebuilt and restarted the LSC model, and confirmed that the anti-whitening filter banks in the PD input filter modules were indeed switching the correct bits.  I'll next put together a script to confirm that the LSC PD whitening is switching as it should.

 

ELOG V3.1.3-